# Metasurface-based Snapshot Shortwave-Infrared Hyperspectral Image Reconstruction with Inter and Intra Prior Learning Network

### v1

@@ II Related work
Apart from Plug-and-play methods, deep unfolding networks unfold deep networks and insert data fidelity term into the training process, thus achieving joint optimization. For instance, ADMM-Net[28] unfolded the alternating direction method of multipliers with a convolutional neural network. DSSP[57] extended the half-quadratic splitting (HQS) method and created a spatial-spectral deep prior to enhance the data fidelity. DGSMP [25] presented an unfolding model estimation framework that utilizes the learned Gaussian scale mixture prior to enhance the model’s performance. RDLUF[36] introduces interaction between different stages with a sequence feature learning formula(similar to GRU[58]), Song et al.[59] implemented feature interaction between different stages using a channel self-attention mechanism, and replaced the process of combining inertia and original gradient terms with an adaptive network.
@@ IV-B Implementation Details
IV-C Comparisons with SOTA Methods

### v2

@@ II Related work
Apart from Plug-and-play methods, deep unfolding networks unfold deep networks and insert data fidelity term into the training process, thus achieving joint optimization. For instance, ADMM-Net[28] unfolded the alternating direction method of multipliers with a convolutional neural network. DSSP[57] extended the half-quadratic splitting (HQS) method and created a spatial-spectral deep prior to enhance the data fidelity. DGSMP [25] presented an unfolding model estimation framework that utilizes the learned Gaussian scale mixture prior to enhance the model’s performance. RDLUF[36] introduces interaction between different stages with a sequence feature learning formula(similar to GRU[58]),
Song et al.[59] implemented feature interaction between different stages using a channel self-attention mechanism, and replaced the process of combining inertia and original gradient terms with an adaptive network.
@@ IV-B Implementation Details
IV-C Evaluation Indexes
We adopt four image quality indexes, including peak signal-to-noise ratio (PSNR) and structure similarity (SSIM)[67], for quantitative evaluation. Specifically, PSNR measures the visual quality, SSIM measures the structure similarity. Generally, bigger values of PSNR and SSIM suggest a better reconstruction accuracy.
IV-D Comparisons with SOTA Methods
@@ IV-E Ablation Study
Ablation of Metasurface Optimization. In order to verify the effectiveness of the our metasurface optimization method combined with the farthest distance sampling and Pearson correlation coefficient, we conduct experiments under the filter array selected by two methods(naive minimize person correlation coefficient method and our method) as shown in table 4. The experimental result indicates that compared with considering the minimum correlation on the whole metasurface database, our proposed method shows great superiority in HSI reconstruction quality. Besides, the visualization results also show that the randomness of the metasurface transmittance curve optimized by our method better than original method, and the correlation between the metasurface units in the filter array changes more smoothly, we believe that the smoothness can improve the stability of the reconstruction network training process, further improve reconstruction performance.